Using the data provided below, please provide: 1. Contribution Margin (CM) 2. Contribution Margin Ratio (CMR) 3. Breakeven in Sales Dollars 4. Margin of Safety (MOS) 5. MOS ratio Units sold: 25,000 Selling price: $ 20 Variable manufacturing costs: $ 155,000 Fixed manufacturing costs: $ 100,000 Variable selling costs: $ 50,000 Fixed administrative costs: $ 30,000.
Answer #1

1. Contribution margin = Sales-Variable manufacturing cost-Variable selling cost = (25000*20)-155000-50000 = 295000

2) Contribution margin ratio = 295000/500000 = 59%

3) Break even sales = Fixed cost/Contribution margin ratio = (100000+30000)/.59 = $220339

4) Margin of safety = (25000*20)-220339 = 279661

5) Margin of safety ratio = 279661/500000 = 55.93%